# Energy-Efficient Balanced Flow Control Achieved through Optimization of Synthetic Jet Placement Based on Deep Reinforcement Learning

### v1

@@ Abstract
This study leverages deep reinforcement learning (DRL) to train synthetic jet-based flow control strategies for circular and square cylinders. The central aim is to ascertain the optimal jet placements that strike an ideal balance between energy efficiency and control effectiveness, by formulating a cohesive strategy based on a comprehensive analysis of flow control performance and energy consumption across a range of configurations. First, the results from single-action training indicate that for circle cylinder, positioning the synthetic jet approximately 105° from the stagnation point achieves the most energy-efficient and effective control strategy. For square cylinder, placing the jet near the rear corner addresses the dual objectives of minimizing energy consumption and maximizing control performance. Second, compared to single-action control, multi-action control exhibits reduced convergence speed and stability. However, simultaneously activating synthetic jets at multiple locations significantly decreases initial energy consumption and enhances energy efficiency. The findings underscore the critical importance of accurately positioning the synthetic jet at the primary flow separation point, as this alignment not only enhances flow control performance but also establishes an optimal balance between energy efficiency and control effectiveness within the flow system. Furthermore, the interaction between synthetic jets and the flow system alters surface streamline patterns and local flow structures, resulting in effects that are considerably larger in scale than the jets themselves. This study crystallizes the potential of DRL algorithms to intelligently optimize intricate flow control strategies, effectively balancing energy efficiency with enhanced flow control performance through strategically optimized jet placement.
@@
Authors should not enter keywords on the manuscript, as these must be chosen by the author during the online submission process.
MSC Codes (Optional) Please enter your MSC Codes here
@@ 1 INTRODUCTION
Active flow control (AFC) involves regulating and controlling fluid flow through external mechanisms such as injection, suction, or vibration to optimize flow characteristics and enhance system efficiency, performance, and stability. Despite significant advancements in this field over the past few decades, there remain unresolved challenges, particularly in addressing the nonlinear complexities inherent in fluid dynamics (Collis & Joslin, 2004; Jahanmiri, 2010). Fluid dynamics problems are intrinsically nonlinear, especially when dealing with turbulent and separated flows, where the flow behavior is extremely complex and difficult to control accurately using traditional analytical or numerical methods (Cattafesta & Sheplak, 2011; Gad-el Hak, 1996). The challenges are further compounded in high-dimensional and multi-scale problems, where finding a global optimum solution is particularly challenging. Consequently, there are still numerous difficulties in optimizing control strategies, achieving real-time control, enhancing system robustness, and reducing energy consumption (Bewley, 2001; Brunton & Noack, 2015a). Addressing these issues will not only help overcome existing technical barriers but also promote the application and development of AFC technologies in a broader range of fields (Gad-el Hak, 1996; Bewley, 2001; Brunton & Noack, 2015b).
@@
The rapid development of artificial intelligence brings new vitality to the field of flow control, with reinforcement learning (RL) emerging as a key machine learning algorithm that effectively addresses complex decision-making and control problems (Brunton et al., 2020; Ren et al., 2020). The core mechanism of RL lies in the agent’s ability to interact with the environment, gradually optimizing its strategy based on rewards or penalties received from its actions, with the goal of maximizing cumulative rewards (Kaelbling et al., 1996; Sutton & Barto, 2018). Reinforcement learning agents possess the ability for long-term planning and decision-making, continuously adjusting their behavior strategies in response to environmental feedback, making it a powerful tool for autonomous learning and decision-making (Arulkumaran et al., 2017). Deep learning (DL) has strong expressive capabilities, allowing it to approximate complex nonlinear functions (LeCun et al., 2015; Janiesch et al., 2021). When RL is combined with DL to form deep reinforcement learning (DRL), this approach leverages the feature extraction capabilities of DL and the strategy optimization strengths of RL (François-Lavet et al., 2018). Consequently, DRL is highly effective at tackling control challenges in high-dimensional, complex state spaces, establishing itself as a powerful technique for advancing flow control technologies and addressing intricate fluid dynamics problems.
The development of DRL traces back to the mid-20th century, with significant progress occurring as deep learning and computational power advance (Kaelbling et al., 1996). A pivotal milestone comes in 2013 when Google DeepMind introduces the Deep Q-Network (DQN), demonstrating DRL’s remarkable potential in mastering Atari games and establishing it as a prominent and rapidly expanding research area (Mnih, 2013; Arulkumaran et al., 2017). Subsequent advancements, particularly in proximal policy optimization (PPO) and distributed DRL techniques, further propel the field. PPO addresses the inherent instability of traditional policy gradient methods by introducing a clipped objective function, ensuring stable and efficient policy updates (Schulman et al., 2017). Distributed DRL, enabling parallel training across multiple machines, greatly improves computational efficiency and scalability (Sutton & Barto, 2018; Liang et al., 2018). These developments lead to DRL applications in various domains, including AlphaGo’s performance in Go, autonomous vehicle navigation, robotic operations, financial trading strategies, and smart manufacturing and healthcare (Li, 2018; Silver et al., 2018). These applications showcase DRL’s exceptional abilities in high-dimensional perception, continuous decision-making, and complex environment handling, highlighting its immense potential and broad impact in practical applications.
@@
Given the exceptional decision-making capabilities of reinforcement learning, its methods and derivative technologies are seeing widespread development in fluid mechanics, covering various areas such as turbulence modeling, flow prediction, shape optimization, and flow control (Rabault et al., 2020; Viquerat et al., 2022; Vignon et al., 2023). In the domain of turbulence modeling, Novati et al. (2021) leveraged multi-agent reinforcement learning to automate the discovery of turbulence models, while Kurz et al. (2023) utilized a multi-agent reinforcement learning framework to identify the optimal eddy viscosity model for implicitly filtered large eddy simulations. In the field of shape optimization, Viquerat et al. (2021) spearheaded the application of DRL for direct shape optimization, demonstrating that artificial neural networks (ANN) could autonomously generate optimal shapes within a constrained time frame and without prior knowledge, given an appropriate reward mechanism. In the context of flow control, Rabault et al. (2019) introduced the first application of an ANN trained via DRL for flow control, successfully reducing drag by 8%. These studies highlight the significant potential and promising prospects of DRL in the field of fluid mechanics.
[…]
Secondly, DRL algorithms have been effectively applied to control complex flow separation phenomena around a wide range of geometries, demonstrating their versatility and robustness. These applications include fluidic pinball systems (Feng et al., 2023), square cylinder (Wang & Xu, 2024d), airfoil (Wang et al., 2022b), elliptical cylinder (Wang & Xu, 2024a), and flat plate (Wang & Xu, 2024a). For instance, Feng et al. (2023) demonstrated that DRL could achieve effective control in fluidic pinball systems, delivering results that were comparable to, and in some cases even surpassed, optimal strategies developed through conventional methods. In another study, Wang & Xu (2024a) applied DRL to active flow control on elliptical cylinders with varying aspect ratios, successfully reducing drag and suppressing vortex shedding. However, they noted that as the aspect ratio decreases, the control challenges increase accordingly. These studies show DRL’s wide applicability and effectiveness in optimizing bluff body flow performance.
Thirdly, significant progress has been made in refining DRL algorithms and optimizing probe placements. Li & Zhang (2022) found that positioning probes in sensitive regions enhances control efficiency, even with fewer probes. Xia et al. (2024) promoted an energy-efficient reinforcement learning approach for drag reduction using partial surface pressure measurements, while Suárez et al. (2024) demonstrated an 8% drag reduction in 3D turbulent flow with DRL-based AFC strategies, achieving better mass cost efficiency than traditional controls. Additionally, Paris et al. (2021a) developed a sensor optimization algorithm that reduced the number of sensors while maintaining high performance. Parallelization has also proven crucial for DRL in flow control. Rabault & Kuhnle (2019) showed that parallelizing simulations accelerates the discovery of control strategies. Wang et al. (2022a) developed DRLinFluids, integrating DRL with OpenFOAM to support further research. Wang & Xu (2024c) improved parallel efficiency, achieving a 47-fold increase in training speed.
[…]
Building on the aforementioned research, we have recognized that the placement of synthetic jets plays a crucial role in determining flow control performance. Moreover, we have observed that many studies lack a thorough evaluation of control costs when assessing flow control effectiveness. For active flow control strategies to be viable, they must achieve the desired outcomes while also ensuring energy efficiency. This necessitates that the control strategy optimizes flow behavior with minimal energy input, thereby avoiding excessive energy consumption due to the control measures themselves. Additionally, the variability of optimal jet positions across different Reynolds numbers has captured our attention. We propose that a fluid dynamics-based analysis could provide unified guidelines for the strategic placement of synthetic jets. To this end, we plan to investigate the placement of synthetic jets for flow control around both square and circular cylinder at $Re=100$.
This study will consider multiple situations, including single-agent control of individual jet pairs, single-agent simultaneous control of multiple jet pairs, and the development of distinct control strategies for each jet pair. Through these configurations, we aim to optimize both control performance and energy efficiency. Crucially, a physical analysis of the training results for flow around circular and square cylinders offers qualitative insights into the design of jet positioning. This analysis is expected to yield a deeper understanding of the underlying physical principles that govern the effectiveness of different jet locations in managing separated flows. The structure of this paper is as follows: § 2 describes the research problem and the framework for coupling DRL with AFC. § 3 and § 4 present the results of the DRL training. In § 5, we analyze and discuss the physical phenomena observed in the training results for both circular and square cylinder flows. Finally, § 6 summarizes the conclusions of this study.
@@ 2 METHODOLOGY
We use DRL algorithms to manipulate synthetic jets for active flow control, focusing on the impact of jet positioning on flow performance around square and circular cylinders at $Re=100$. The fundamental numerical methods, software, and key parameters are outlined in § 2.1. In § 2.2, we describe the flow characteristics around a circular cylinder and the arrangement of the synthetic jets, while § 2.3 provides corresponding details for a square cylinder. § 2.4 introduces the core concepts, mathematical models, and distinctions of deep learning and reinforcement learning, with a detailed discussion of the PPO algorithm. Lastly, § 2.5 demonstrates the application of DRL algorithms to address flow control challenges, offering a comprehensive overview of the agent-environment framework and related key concepts.
@@ 2.1 Numerical simulation
We describe the behavior of incompressible viscous fluid flow within the domain $\Omega\subset\mathbb{R}^{nd}$ over the time interval $(0,T)$ using the Navier-Stokes equations. These equations govern the evolution of the fluid velocity field $\mathbf{u}=\mathbf{u}(\mathbf{x},t)$ and pressure field $p=p(\mathbf{x},t)$, where $\mathbf{x}$ represents the spatial coordinates and $t$ denotes time. The governing equations in their non-dimensional form are expressed as:
@@
where $\mathbf{u}$ denotes the non-dimensional velocity, $t$ represents the non-dimensional time, and $p$ is the non-dimensional pressure. The Reynolds number is defined as $Re=\frac{UD}{\nu}$, where $\nu$ denotes the kinematic viscosity of the fluid.
This study employs OpenFOAM for numerical simulations, a widely used and flexible Navier-Stokes solver that is highly regarded in both industry and academia (Jasak et al., 2007). In the process of solving partial differential equations using the finite volume method, we divide the computational domain into a finite number of control volumes to more accurately capture and analyze the complex behavior of fluids. The pimpleFoam solver is specifically designed for addressing unsteady Navier-Stokes equations. It integrates the Semi-Implicit Method for Pressure-Linked Equations (SIMPLE) with the Pressure-Implicit with Splitting of Operators (PISO) algorithms (Issa, 1986; Jang et al., 1986). This integration enables the effective resolution of complex, time-dependent flow scenarios, allowing for larger time steps while maintaining both stability and computational efficiency.
@@ 2.2 Circular cylinder environment
Figure 1: Numerical simulation model of flow around a cylinder. (a) Domain configuration and boundary conditions. (b) Details of the jets conditions on the cylinder. Jets are located at $\Omega=90^{\circ}$ and $\Omega=270^{\circ}$ on the cylinder, with a jet width of $\omega=10^{\circ}$. Parabolic velocity distribution is used for each jet. (c) Mesh structure around the object, and (d) Zoomed-in view of the mesh near the cylinder boundary.
The configuration employed in this study for simulating flow around a circular cylinder closely follows the classical benchmark established by Schäfer et al. As illustrated in figure 1, the cylinder is placed within a rectangular computational domain with dimensions of $22D$ along the $x$-axis and $4.1D$ along the $y$-axis. The origin of the Cartesian coordinate system is positioned at the center of the cylinder. The cylinder is slightly offset in the $y$-direction to induce vortex shedding, facilitating the analysis of oscillatory flow phenomena. The boundaries of the computational domain are divided into an inlet $\Gamma_{\text{in}}$, an outlet $\Gamma_{\text{out}}$, no-slip walls $\Gamma_{\text{w}}$, and two separate jets on the cylinder $\Gamma_{i}\,(i=1,2)$, as shown in figure 1(a). At the inlet $\Gamma_{\text{in}}$, the inflow velocity along $x$-axis is prescribed by a parabolic velocity profile in the form,
$U_{\text{inlet}}(y)=U_{m}\frac{(H-2y)(H+2y)}{H^{2}},$ (2)
and that along $y$-axis is prescribed as,
$V_{\text{inlet}}(y)=0.$ (3)
$U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain. The average inlet velocity $\overline{U}$, is related to the parabolic velocity profile $U_{\text{inlet}}(y)$ through the expression:
@@
Flow control on the cylinder surface is achieved by introducing synthetic jets at the top and bottom points, as shown in figure 1(b). These jets, denoted as $\Gamma_{i}$ $(i=1,2)$, can inject or suction fluid to control the flow around the cylinder. Each jet spans an angular width of $\omega=10^{\circ}$ and is oriented perpendicularly to the outer surface of the cylinder. The velocity profile within each jet follows a parabolic distribution, allowing for both positive (blowing) and negative (suction) velocities. A key aspect of the AFC implementation is ensuring mass conservation within the system. This is achieved by balancing the net mass flow rates of the jets, i.e., $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, which ensures that the net flow into or out of the system remains zero.
Additionally, a no-slip boundary condition is enforced on the areas of the cylinder surface where the jets are not located, preventing any relative motion between the fluid and the cylinder in those regions. At the outlet boundary $\Gamma_{\text{out}}$, an outflow condition is applied, where the velocity is extrapolated based on the internal flow field. No-slip boundary conditions are also applied at the upper and lower walls $\Gamma_{\text{w}}$, ensuring that the velocity at these walls is zero relative to the boundary surface. This specification of boundary conditions ensures accurate flow simulation and effective control through synthetic jets. Mathematically, the boundary conditions are written as:
@@
where, $\mathbf{n}$ denotes the normal vector, $U$ is the inflow velocity, and $f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets.

### v5

# Strategies for energy-efficient flow control leveraging deep reinforcement learning

@@ Abstract
This study investigates active flow control in two-dimensional flows at a Reynolds number of 100 using Deep Reinforcement Learning (DRL). We utilize DRL to develop flow control strategies that enhance energy efficiency and minimize energy consumption, thereby addressing the limitations of traditional methods. We find that the optimal jet placement for both square and circular cylinders is at the main flow separation point, achieving the best balance between energy efficiency and control effectiveness. For the circular cylinder, positioning the jet at approximately 105° from the stagnation point requires only 1% of the inlet flow rate and achieves an 8% reduction in drag, with energy consumption one-third of that at other positions. For the square cylinder, placing the jet near the rear corner requires only 2% of the inlet flow rate, achieving a maximum drag reduction of 14.4%, whereas energy consumption near the front corner is 27 times higher, resulting in only 12% drag reduction. In multi-action control, the convergence speed and stability are lower compared to single-action control, but activating multiple jets significantly reduces initial energy consumption and improves energy efficiency. Physically, the interaction of the synthetic jet with the flow generates new vortices that modify the local flow structure, significantly enhancing the cylinder’s aerodynamic performance. Our control strategy achieves a superior balance between energy efficiency and control performance compared to previous studies, underscoring its significant potential to advance sustainable and effective flow control.
@@
Flow control, deep reinforcement learning, drag reduction, vortex shedding
[…]
organization=State Key Laboratory of Ocean Engineering, School of Ocean and Civil Engineering, Shanghai Jiao Tong University,city=Shanghai, postcode=200240, country=China
@@ 1 INTRODUCTION
Active flow control (AFC) involves regulating and controlling fluid flow through external mechanisms such as injection, suction, or vibration to optimize flow characteristics and enhance system efficiency, performance, and stability. Despite significant advancements in this field over the past few decades, there remain unresolved challenges, particularly in addressing the nonlinear complexities inherent in fluid dynamics Collis and Joslin (2004). Fluid dynamics problems are intrinsically nonlinear, especially when dealing with turbulent and separated flows, where the flow behavior is extremely complex and difficult to control accurately using traditional analytical or numerical methods Cattafesta and Sheplak (2011). These challenges are intensified in high-dimensional, multi-scale problems, making global optimization, real-time control, robustness, and energy efficiency difficult to achieve Bewley (2001).
@@
The rapid development of artificial intelligence has brought new vitality to the field of flow control. Reinforcement learning (RL) has emerged as a key machine learning algorithm that effectively addresses complex decision-making and control problems Brunton et al. (2020); Brunton and Noack (2015). RL agents continuously adjust their behavior strategies in response to environmental feedback, making RL a powerful tool for autonomous learning and decision-making Arulkumaran et al. (2017). Deep learning (DL) has strong expressive capabilities, allowing it to approximate complex nonlinear functions Janiesch et al. (2021). When combined with RL to form deep reinforcement learning (DRL), this approach leverages DL’s feature extraction capabilities and RL’s strategy optimization strengths François-Lavet et al. (2018). DRL originated in the mid-20th century, with deep learning and computational power driving its development. The 2013 introduction of Deep Q-Network by Google DeepMind was a major breakthrough, catalyzing rapid growth Mnih (2013). Innovations like PPO have further enhanced DRL’s stability and efficiency, enabling its application in fields such as autonomous navigation and manufacturing Li (2018); Schulman et al. (2017).
@@
Given the exceptional decision-making capabilities of reinforcement learning, its methods and derivative technologies are seeing widespread development in fluid mechanics, covering various areas such as turbulence modeling, flow prediction, shape optimization, and flow control Rabault et al. (2020); Viquerat et al. (2022); Vignon et al. (2023). In the domain of turbulence modeling, Novati et al. (2021) leveraged multi-agent reinforcement learning to automate the discovery of turbulence models, while Kurz et al. (2023) utilized a multi-agent reinforcement learning framework to identify the optimal eddy viscosity model for implicitly filtered large eddy simulations. In the field of shape optimization, Viquerat et al. (2021) spearheaded the application of DRL for direct shape optimization, demonstrating that artificial neural networks (ANN) could autonomously generate optimal shapes within a constrained time frame and without prior knowledge, given an appropriate reward mechanism. In the context of flow control, Rabault et al. introduced the first application of an ANN trained via DRL for flow control, successfully reducing drag by 8%.
AFC techniques combining computational fluid dynamics (CFD) with DRL have been extensively studied for robustness across Reynolds numbers ($Re$), flow separation control, parallelization, algorithm comparisons, and probe optimization. Firstly, DRL controllers have shown strong generalization across various $Re$. Tang et al. achieved significant drag and lift reductions in cylinder flows, while Jia and Xu demonstrated the DRL agent’s ability to fully suppress vortex shedding across multiple $Re$. Ren et al. and Wang et al. further explored DRL in higher $Re$ and 3D flows. Secondly, DRL algorithms have been successfully applied to control complex flow separation phenomena across various configurations, including fluidic pinball systems Feng et al. (2023), square cylinders Jia and Xu (2024), airfoils Wang et al. (2022), elliptical cylinders Jia and Xu (2024), and flat plates Jia and Xu (2024), showcasing their versatility and robustness. Thirdly, significant advancements have been achieved in refining DRL algorithms and optimizing probe placements, as highlighted by Li and Zhang; Xia et al.; Suárez et al.; Paris et al. Additionally, parallelization has proven critical for enhancing the efficiency of DRL-based flow control, with studies such as Rabault and Kuhnle, Jia and Xu and Wang et al. demonstrating its potential to accelerate training and strategy discovery.
Despite prior studies exploring DRL-based AFC from various perspectives, research in the field of synthetic jet technology has identified that the placement of jet actuators plays a crucial role in determining control effectiveness and energy consumption Zhu et al. (2019); Zong et al. (2018). The strategic layout of synthetic jets plays a pivotal role in determining their interaction with boundary layer flows, thereby influencing both control performance and energy efficiency Lin (2002); Qayoum et al. (2010); Ho et al. (2024). As a highly effective flow control method, the optimal positioning of synthetic jets directly dictates their interaction with the boundary layer and the main flow, affecting aerodynamic properties such as drag reduction and lift modulation Chen et al. (2022); Ho et al. (2024); Girfoglio et al. (2015). Moreover, different jet locations lead to varying levels of energy utilization efficiency, where improper placement may result in increased energy consumption with limited control benefits Saidur et al. (2010); Zuckerman and Lior (2006). Investigating the coupling between synthetic jet placement and control performance is essential for understanding mechanisms, yet systematic evaluation of this balance remains limited.
Recent studies have begun to explore the influence of synthetic jet actuator placement on flow control. Chen et al. demonstrated that placing actuators near the rear corners of a square cylinder at $Re=100$ improved vortex shedding suppression. At higher $Re$, Yan et al. showed that positioning actuators near the front corners enhanced performance at $Re=2000$. However, two key issues remain: firstly, these studies primarily focus on control effectiveness without providing a detailed analysis of the underlying physical mechanisms. Secondly, their qualitative findings on optimal jet placement vary across different $Re$, leading to inconsistent results Chen et al. (2023); Yan et al. (2023). This has motivated research exploring the impact of synthetic jet location and width on flow control performance at $Re=100$ and 500, utilizing an entropy-regularized reinforcement learning approach Jia and Xu (2024). Despite these advances, there remains a lack of comprehensive research on the sensitivity of synthetic jet placement to flow separation suppression in bluff body, as well as fluid mechanics-based qualitative coupled with quantitative analysis.
We note that Yan et al. highlighted the importance of multiple jet actuators, where configurations with multiple independent actuators demonstrated optimal performance. Thus, we have developed a strong interest in comparing the control performance of multiple synthetic jets operating simultaneously with that of single jets working independently. Firstly, the synergy among multiple jets can induce nonlinear flow behaviors, resulting in control effects that differ significantly from those achieved by a single actuator. Evaluating these performance differences can reveal the potential and limitations of synthetic jets in flow control. Secondly, as the number of actuators increases, the interactions between jets and the complexity of energy distribution also rise, potentially impacting energy consumption and control efficiency. A systematic comparison of multiple versus single-pair synthetic jet activation offers insights into achieving efficient, energy-saving flow control.
Given the identified challenges and research gaps in multiple aspects, we conduct a comprehensive analysis of synthetic jet placement to evaluate its impact on flow control performance and energy efficiency. We have designed five synthetic jet placements for cylindrical bodies, resulting in a total of nine jet configuration schemes, including five control strategies for single synthetic jet actions and four for simultaneous multi-jet actions. Additionally, for square cylinders, we designed four jet placements, leading to seven jet configuration schemes in total, comprising four control strategies for single synthetic jet actions and three for simultaneous multi-jet actions. The innovative contributions are summarized as follows:
  • 1.
    Sensitivity analysis of synthetic jet placement: Studies by Tang et al. and Rabault et al. highlight DRL’s superior decision-making in effective flow control. Tang et al. achieved a 5.7% reduction in drag at $Re=100$ using two pairs of synthetic jets. In contrast, Rabault et al., employing the same control algorithm under similar baseline flow conditions, achieved an 8% drag reduction with only a pair of synthetic jets. These findings underscore the critical impact of synthetic jet number and placement on control performance, warranting further investigation. Building on previous studies Tang et al. (2020); Rabault et al. (2019); Yan et al. (2024); Suzuki (2006); Ma et al. (2024), we recognize that the placement of synthetic jets plays a pivotal role in determining control performance, directly influencing the effectiveness and energy efficiency of flow control strategies.
  • 2.
    Comparison of single-jet and multi-jet control: Yan et al. demonstrated in their study of flow around a square cylinder that multiple synthetic jet actuators outperform a single pair, highlighting the potential advantages of multi-jet configurations. However, this conclusion contrasts with findings from studies on circular cylinder flows Tang et al. (2020); Rabault et al. (2019). To address this discrepancy, our study uniquely compares the performance of single-jet and multi-jet actuation in both square and circular cylinder flow scenarios. By exploring the synergistic effects of multi-jet actuation, we aim to uncover nonlinear flow interactions and provide insights into optimizing actuator design for balanced control efficiency and energy consumption.
  • 3.
    Achieving complete suppression of vortex shedding: previous studies have achieved notable success in drag reduction; however, a common issue persists: the controlled flow in their results still exhibit vortex shedding, leading to oscillations in both lift and drag coefficients, even at $Re=100$ Rabault et al. (2019); Tang et al. (2020); Paris et al. (2021); Mao et al. (2022); He et al. (2023). This observation prompted us to delve deeper into the underlying mechanisms. Drawing on vortex-induced vibration theory Sarpkaya (1979), it is reasonable to assume that the residual oscillations in lift and drag coefficients are directly linked to the incomplete suppression of vortex shedding in the controlled flow. We tackle this challenge with a DRL-based active flow control strategy that completely eliminates vortex shedding and fully stabilizes the flow. Unlike previous studies reporting persistent instabilities, our results show lift and drag coefficients without any oscillatory behavior under identical conditions.
  • 4.
    Enhancing energy efficiency in flow control: another significant contribution of this study is the exploration of energy-efficient control strategies. We observe that existing studies fail to comprehensively evaluate the control costs when assessing flow control performance. For active flow control strategies to be feasible, they must not only achieve the desired outcomes but also ensure energy efficiency. This requires the control strategies to optimize flow behavior with minimal energy input, thus preventing excessive energy consumption due to the control measures themselves. We not only ensures stable performance but also guarantees that the control solutions are optimized for minimal energy consumption.
The structure of this paper is organized as follows: In section 2, we introduce the research problem and detail the framework for integrating DRL with AFC. Section 3 focuses on the DRL training results for circular cylinders, highlighting the optimization of jet placement and control strategies. Similarly, section 4 presents the results for square cylinders, emphasizing the unique flow characteristics and control challenges. In section 5, we analyze the physical phenomena underlying the DRL-trained control strategies for both cylinder geometries, offering insights into the flow dynamics and the effects of jet configurations. Finally, section 6 summarizes the main findings of this study, discusses the implications for actuator design.
@@ 2 METHODOLOGY
DRL algorithms are employed to manipulate synthetic jets for active flow control, focusing on the effect of jet positioning on flow performance around square and circular cylinder at $Re=100$. The numerical methods, software, and key parameters are outlined in section 2.1. Section 2.2 describes the flow characteristics and jet arrangements for circular cylinders, while section 2.3 provides similar details for square cylinders. Section 2.4 introduces the core concepts, mathematical models, and distinctions of DRL, with an emphasis on the PPO algorithm. Finally, section 2.5 demonstrates the application of DRL algorithms to address flow control challenges.
@@ 2.1 Numerical simulation
The behavior of incompressible viscous fluid flow within the domain $\Omega\subset\mathbb{R}^{nd}$ over the time interval $(0,T)$ is described using the Navier-Stokes equations. These equations govern the evolution of the fluid velocity field $\mathbf{u}=\mathbf{u}(\mathbf{x},t)$ and pressure field $p=p(\mathbf{x},t)$, where $\mathbf{x}$ represents the spatial coordinates and $t$ denotes time. The governing equations in their non-dimensional form are expressed as:
@@
where the Reynolds number is defined as $Re=\frac{UD}{\nu}$, with $U$ being the upstream inlet velocity, $D$ as the characteristic length (the cylinder diameter for a circular cylinder and the length of one side for a square cylinder), and $\nu$ denoting the kinematic viscosity of the fluid.
We employ OpenFOAM, a widely used Navier-Stokes solver recognized in both industry and academia Jasak et al. (2007), for numerical simulations. The computational domain is discretized using the finite volume method to capture complex fluid behavior. The pimpleFoam solver, combining the SIMPLE and PISO algorithms Issa (1986); Jang et al. (1986), efficiently resolves unsteady Navier-Stokes equations, enabling larger time steps with stability and computational efficiency.
@@ 2.2 Circular cylinder environment
Figure 1: Schematics of the physical model of flow around a circle cylinder. (a) Domain configuration and boundary conditions. (b) Details of the jets conditions on the cylinder. (c) Mesh structure around the cylinder, and (d) Zoomed-in view of the mesh near the cylinder boundary.
@@
$U_{\text{inlet}}(y)=U_{m}\left((H-2y)(H+2y)/H^{2}\right),$ (2)
and that along $y$-axis is prescribed as $V_{\text{inlet}}(y)=0$. $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain. The average inlet velocity $\overline{U}$, is related to the parabolic velocity profile $U_{\text{inlet}}(y)$ through the expression:
[…]
Additionally, a no-slip boundary condition is enforced on the areas of the cylinder surface where the jets are not located. At the outlet boundary $\Gamma_{\text{out}}$, an outflow condition is applied, where the velocity is extrapolated based on the internal flow field. No-slip boundary conditions are also applied at the upper and lower walls $\Gamma_{\text{w}}$. This specification of boundary conditions ensures accurate flow simulation and effective control through synthetic jets. Mathematically, the boundary conditions are written as:
@@
where $\mathbf{n}$ denotes the normal vector and $f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets. Flow control is implemented through synthetic jet pairs positioned at the cylinder, as illustrated in figure 1(b). These synthetic jet pairs, denoted as $\Gamma_{i}$ $(i=1,2)$, alternately inject and suction fluid to actively manipulate the surrounding flow field.